OpenStudy (superhelp101):

Solve the following system of equations. 2x - 3y + z = -7 3x + y - 2z = 13 x - 2y + 3z = -10 (-2, 1, 0) (–2, 3, 2) (2, 9, 2) (2, 3, -2)

OpenStudy (superhelp101):

I got D. Just need to check. :)

OpenStudy (superhelp101):

@tkhunny @Whitemonsterbunny17 @aum @zzr0ck3r @mathstudent55 @dumbcow anyone? :)

OpenStudy (aum):

To check if your answer is correct, plug the answer into the original equations and see if they are valid.

OpenStudy (superhelp101):

ok :)

OpenStudy (superhelp101):

idk I got something different

OpenStudy (aum):

Your solution: x = 2, y = 3, z = -2 2x - 3y + z = -7 2(2) - 3(3) - 2 4 - 9 - 2 -7 so this is true. Check the same way with the other two equations.

OpenStudy (superhelp101):

I think i got the same thing.

OpenStudy (superhelp101):

can you check the other two also?

OpenStudy (aum):

3x + y - 2z = 13 3(2) + 3 - 2(-2) 6 + 3 + 4 13 (true)

OpenStudy (aum):

x - 2y + 3z = -10 2 - 2(3) + 3(-2) 2 - 6 - 6 -10 (true)

OpenStudy (superhelp101):

thank you so much!

OpenStudy (aum):

you are welcome.